OpenStudy (anonymous):

find the difference: 4/x-5/x+5

OpenStudy (anonymous):

With no parentheses and using normal precedence rules:\[4/x-5/x+5 \]\[\frac{4}{x}-5/x+5 \]\[\frac{4}{x}-\frac{5}{x}+5 \]\[-\frac{1}{x}+5 \]\[5-\frac{1}{x} \]